OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (anonymous):

Let x = # of boys y = # of girls So... x+y=94 x=2y+22 Now, you can use substitution to figure this out. x+y=94 2y+22+y = 94 Once you find out the number of girls, sub it back into the second equation to get the number of boys